Please help me solve this I think of a number .I subtract 14 from the number and then treble the difference .The result is two-third of the number .Find the original number
Answer by swincher4391(1107)

3*(N-14) = 2/3*N
3N - 42 = 2/3* N
9/3N - 42 = 2/3N
7/3N = 42
N = 42(3/7)
N = 18
Check:
18-14 = 4
4*3 = 12
18(2/3) = 12 CHECK!